OpenStudy (anonymous):

Find two positive numbers that satisfy the given requirements. The second number is the reciprocal of the first and the sum is a minimum.

OpenStudy (anonymous):

minimize \[x+\frac{1}{x}\]

OpenStudy (anonymous):

take the derivative and find the critical points

OpenStudy (anonymous):

\[y=x+\frac{1}{x}\] \[y'=1-\frac{1}{x^2}\] set \[1-\frac{1}{x^2}=0\] get \[1=\frac{1}{x^2}\] so \[x=\pm1\] by inspection. since you want a positive number the answer is 1

OpenStudy (anonymous):

which frankly you could have arrived at by ignoring calc and just thinking